### v3

# Network Formation and Dynamics Among Multi-LLMs

Abstract.
Social networks shape opinions, behaviors, and information dissemination in human societies. As large language models (LLMs) increasingly integrate into social and professional environments, understanding their behavior within the context of social interactions and networks becomes essential. Our study analyzes LLMs’ network formation behavior to examine whether the dynamics of multiple LLMs are similar to or different from human social dynamics. We observe that LLMs exhibit key social network principles, including preferential attachment, triadic closure, homophily, community structure, and the small-world phenomenon, when asked about their preferences in network formation. We also investigate LLMs’ decision-making based on real-world networks, revealing that triadic closure and homophily have a stronger influence than preferential attachment and that LLMs perform well in network formation predictions. Overall, our study opens up new possibilities for using LLMs in network science research and helps develop socially aware LLMs by shedding light on their social interaction behaviors and exploring their impacts on social dynamics.
large language models, networks, network formation, link prediction, multi-agent systems
conference: ; ;
Introduction
Recent progress in large language models (LLMs), such as GPT (OpenAI, 2023) and Llama 2 (Touvron et al., 2023), have shown promising developments in AI techniques and their integration into real-life applications. It is thus crucial to comprehend AI actions to ensure they align with human expectations, mitigate potential risks, and maximize their benefits. Misaligned AI actions may lead to unintended consequences, such as biased decision-making, fairness issues, and the miscoordinative or non-cooperative behavior (Rahwan et al., 2019). Recently, researchers have started to apply social science methodologies, such as methods analogous to laboratory experiments (Horton, 2023; Aher et al., 2023; Veselovsky et al., 2023; Manning et al., 2024), agent-based modeling (Park et al., 2018; Gao et al., 2023; He et al., 2023; De Marzo et al., 2023; Fatemi et al., 2023; Perozzi et al., 2024), and qualitative methods, to study LLMs. These methods not only reveal the capabilities and interpretability of LLMs but also suggest their potential for applications in social science (Horton, 2023; Park et al., 2023; Chen et al., 2023; Leng et al., 2023).
In human societies, social networks play a crucial role in shaping individual behaviors, preferences, and connections, as well as influencing the diffusion of information and norms across communities (Rogers, 2003; Bakshy et al., 2012; Fowler and Christakis, 2008; Banerjee et al., 2013; Yuan et al., 2018). LLMs have shown great potential in social contexts, notably as intelligent personal assistants that facilitate social and prosocial interactions (see, e.g., (Papachristou et al., 2023; Chopra and Haaland, 2023; Veselovsky et al., 2023)). However, less is known about how LLMs’ innate behaviors and preferences align with human network formation principles (Jackson, 2008; Zhou et al., 2023; Park et al., 2023). This is particularly crucial, as it sheds light on the potential of these models to shape and be shaped by the networks of human relationships, which is a fundamental aspect of social systems.
Our study thus explores LLMs’ behaviors and preferences in the context of network formation with both synthetic and real-world social networks. By analyzing interactions between multiple LLMs (or multi-LLMs), we aim to understand the implications of LLMs representing humans in social and professional settings. Specifically, we examine micro-level social network properties including preferential attachment (Barabási and Albert, 1999), triadic closure (Granovetter, 1973), and homophily (McPherson et al., 2001), as well as macro-level properties including community structure (Newman and Girvan, 2004), and the small-world phenomenon (Kleinberg, 2000; Watts and Strogatz, 1998).
In simulations conducted using synthetic networks, we have the following findings: First, we observe that LLMs can generate networks exhibiting preferential attachment and following scale-free degree distributions. Moreover, LLMs demonstrate strong tendencies toward triadic closure and exhibit tendency of homophily. These tendencies further explain the emergence of community structures in networks formed by multi-LLMs. Additionally, we show that LLM agents exhibit dynamics consistent with small-world phenomena akin to those observed in the Watts-Strogatz model (Watts and Strogatz, 1998).
Next, we compare between these principles to identify the key factors driving network dynamics in real-world network formation contexts. To achieve this, we consider real-world networks and ask each LLM to simulate a node in the network. We provide information on their network structure and ask for their decisions for the next link formation or dissolution step. We approach this problem as a discrete choice modeling problem (Liben-Nowell and Kleinberg, 2003; Overgoor et al., 2019). In this real-world setting, we find that LLMs’ choices are primarily weighted on homophily and triadic closure, compared to preferential attachment, and that LLMs substantially exceed random guessing when predicting new links. Additionally, the resulting choices of LLM agents strengthen the network’s community structure.
Our study contributes to the literature on LLMs in multiple ways. First, it advances the discussion on aligning LLMs with human preferences, which has potential applications in developing socially aware LLMs for real-life interactions. Insights into LLMs’ innate behaviors in network formation can inform the design of LLMs for effective collaboration in social and professional settings. Second, our study could contribute to understanding the impact of LLMs on social dynamics and norms, such as their role in shaping online communities’ interaction patterns or influencing the adoption of new communication norms in professional settings. Finally, our work has direct applications to synthetic dataset generation, a problem important for the network science community. By controlling for a few parameters, such as the temperature, our work can be utilized to generate a variety of datasets, as well as generate synthetic datasets from real-world data to protect privacy.
(a) Structure-based
(b) Degree-based
Figure 1. Results for Principle 1 (preferential attachment): We display simulated networks comprising 200 nodes across different temperatures. For structure-based simulations (Figure 1(a)), the multi-LLM setup was given neighborhood information $\{N_{j,t}:j\in V_{t}\}$, while for degree-based simulations (Figure 1(b)), node degree data $\{d_{j,t}:j\in V_{t}\}$ was provided ($V_{t}$ corresponds to the vertex set of the network $G_{t}$ at round $t$). The degree distributions were then compared with those of a BA network with $n=200$ nodes. (a): Power-law degree distributions are evident ($P>0.5$, K-S test), with the networks at a temperature of 1.5 closely resembling the BA model ($P>0.1$, K-S test). (b): With degree information only, the networks form more unrealistic star-like structures, diverging from scale-free configurations and more closely mirroring a core-periphery network structure.
Results
In our study, we leveraged state-of-the-art LLMs through the OpenAI Application Programming Interface (API). Our methodology involved constructing multiple LLMs, each acting as an independent agent within separate conversational threads. Each LLM agent was prompted with network information about its peers and faced with decision-making scenarios. For synthetic network simulations, we employed GPT-3.5 agents, whereas GPT-4 agents were used for real-world network analysis. [^1] We utilized the gpt-3.5-turbo model for GPT-3.5 and the gpt-4-1106-preview model for GPT-4, both provided by OpenAI. Notably, both GPT-3.5 and GPT-4 yielded consistent results.
We adopt a multi-LLM framework, emphasizing the multi-agent aspect of our experimental design, which shares similar spirits with recent research in the field (Leng et al., 2023; Akata et al., 2023; Brookins and DeBacker, 2023; Aher et al., 2023; De Marzo et al., 2023). To assess the robustness of our findings, we varied the ‘temperature’ settings in LLMs, which modulate the output variability, in a dynamic network environment. Additionally, we adopted an interview-like method to probe the rationale behind the LLMs’ decision-making processes, enriching our understanding of their behavior in network simulations.
In our analysis, we begin with an examination of three fundamental micro-level network principles: preferential attachment, triadic closure, and homophily. Subsequently, we shift our focus to two macro-level network phenomena: community structure and the small-world phenomenon. Finally, we assess the resilience of these principles as LLM agents interact with real-world networks. This allows us to compare and contrast the network formation preferences between LLMs and humans.
Micro-Level Properties
Principle 1: Preferential Attachment
Preferential attachment is a fundamental concept in network science, illustrating how nodes in a network gain connections over time, leading to a scale-free degree distribution characterized by a few highly connected nodes (Barabási and Albert, 1999; Bianconi and Barabási, 2001).
In our study, we investigate whether LLMs manifest the preferential attachment mechanism. Starting with an empty network, we introduce nodes sequentially at each time step $t$. A new node $i_{t}$ is added during time step $t$, with the network’s existing nodes providing context about their connections. We generated networks with $n=200$ nodes, a size large enough to observe meaningful patterns in the degree distribution. Our simulations include two distinct types: structure-based, where the network’s structure was provided, and degree-based, where only the degree information of potential connections was given. [^2] For details on how features were represented in the prompts, refer to SI Section 5.
The graphical outputs from our structure-based simulation, as illustrated in Figure 1(a), reveal a distinct pattern where a small number of nodes possess a high number of connections, whereas the majority exhibit fewer links. This pattern suggests a scale-free degree distribution, characterized by the probability $\pi(d)$ that a node has degree $d$, mathematically represented as:
(1) $\pi(d)\propto d^{-\gamma},\quad\text{where }\gamma>1.$
Interestingly, we also observe that the number of highly connected “super nodes” increases with temperature.
[…]
Furthermore, in degree-based simulations in a similar spirit of ((De Marzo et al., 2023); see Figure 1(b)), we noted formations resembling a “star” around central nodes. These distributions deviate from the scale-free patterns seen in statistical models like the BA model ($P<0.001$, two-sample K-S test), aligning more closely with core-periphery structures (Borgatti and Everett, 2000).
To interpret the LLM agents’ decision-making, we analyzed the reasoning they provided, focusing on the structure-based results (Figure 1(a)) [^4] The degree-based results’ reasoning is not included as decisions were based solely on degree information.. Two primary reasons emerged, varying with temperature (detailed in SI Section 3):
  • Well-Connectedness: Relating to the node’s degree (number of connections).
  • Mutual Friends: Relating to the number of shared connections.
Interestingly, as the temperature increased, the prevalence of the well-connectedness rationale decreased, while the mutual friends’ rationale increased.
In conclusion, our analysis demonstrates that LLMs, when provided with network structure, can replicate the preferential attachment phenomenon.
Principle 2: Triadic Closure
(a) Structure-Based
(b) Common-Neighbor-Based
Figure 2. Results for Principle 2 (triadic closure). (a) Figure 2(a) shows the resulting networks created by triadic closure when the intersection of the neighborhoods of the query node and each alternative is provided. The node colors correspond to the stochastic block model (SBM) groups to which each node belongs. The red edges correspond to the newly created edges between nodes of the red cluster, the blue edges correspond to the newly created edges between nodes of the blue cluster, and the orange edges correspond to edges created between the two clusters. The newly formed edges are drawn with a larger line width. We observe that the probability of forming an edge within the same community and the marginal transitivity, which indicate triadic closure, is significantly larger than randomly creating links ($P<0.001$, t-test). (b) Figure 2(b) shows the same networks with the only change that instead of the intersection of neighborhoods between the query node and each alternative, we provide the number of common neighbors (i.e., the size of the intersection) between the query node and each alternative. Similarly, we observe that the probability of forming an edge within the same community and the marginal transitivity, which indicate triadic closure, is significantly larger than randomly creating links ($P<0.001$, t-test). The error bars correspond to 95% confidence intervals.
The second micro-level principle we examine is triadic closure, which posits that individuals are more likely to form connections with friends of friends, thus creating closed triads in the network. This process strengthens network structure and cohesion, grounded in the idea that two nodes are more likely to connect if they share a common neighbor (Granovetter, 1973; Mosleh et al., 2024).
To investigate triadic closure, we employ an assortative stochastic block model (SBM) (Newman, 2003) to create an initial network $G_{1}$ with $n$ nodes divided into two equal-sized clusters $A$ and $B$. Connections within each cluster are formed with a probability of $0.5$, while inter-cluster connections occur with a probability of $0.1$. This setup mirrors our assumption that nodes within the same cluster are more inclined to connect due to a higher number of shared neighbors. In subsequent time steps, we then examine each node $i$, considering the intersection of neighborhoods of $i$’s non-neighbors or the number of mutual friends between $i$ and any non-neighbor $j$ [^5] It’s worth noting that similar outcomes arise when providing neighbors instead of common neighbors.
We conducted ten simulations with $n=50$ nodes to facilitate clear visualization and ensure statistical significance [^6] Choosing $n=50$ instead of a larger number like $n=200$ aids in visualization and maintains statistical significance.. To contrast LLM-based network formation, we refer to a random null model, selecting nodes at random for connection.
For evaluating triadic closure, we utilize two metrics: marginal transitivity and probability of edge formation within the same community. Marginal transitivity ($D$) represents the change in the ratio of closed triangles to all triads, transitioning from the initial network $G_{T}$ to the SBM-generated $G_{1}$:
$D=3\times\frac{\text{\# triangles}(G_{T})}{\text{\# triads}(G_{T})}-3\times\frac{\text{\# triangles}(G_{1})}{\text{\# triads}(G_{1})}.$
where a large positive $D$ indicates a strong triadic closure tendency.
The probability of forming an edge within the same community ($\hat{p}$) is calculated by the ratio of edges in $G_{T}\setminus G_{1}$ (newly formed edges) connecting nodes within the same cluster:
$\hat{p}=\frac{\left|\left\{\{i,j\}\in E(G_{T})\setminus E(G_{1}):y_{i}=y_{j}\right\}\right|}{\left|E(G_{T})\setminus E(G_{1})\right|},$
where $y_{i},y_{j}\in{A,B}$ denote the community memberships of nodes $i$ and $j$, respectively. A value of $\hat{p}$ exceeding $0.5$ suggests a triadic closure tendency. As we investigate under SBM, same community membership indicates more open triads being closed.
In Figure 2(a), sample networks are displayed, with the upper panel showing networks where the entire structure is provided and the lower panel showing those with only common neighbor numbers provided. Nodes are color-coded to indicate their cluster memberships in the SBM, with red and blue edges within clusters and orange edges between clusters. Newly formed edges are highlighted with thicker lines.
Marginal transitivity, presented in the fourth panel of each row, shows a statistically significant increase across all temperatures ($P<0.001$, t-test), and compared to the Random model, the increase is notable ($P<0.001$, t-test).
Regarding the probability of within-community network formation (last panels in each row), we found that across all temperatures, this probability significantly exceeds that in the random network formation scenario ($P<0.001$, t-test against Random), indicating an LLM tendency towards triadic closure.
These findings affirm that LLM agents tend to follow triadic closure. We investigate the reasons provided by LLMs for their decisions in SI Section 3.777Reasons for decisions based on the number of common neighbors are not provided, as these decisions are solely based on this information. The analysis revealed that many decisions were influenced by the number of mutual friends.
Principle 3: Homophily
Homophily reflects the tendency for nodes with similar characteristics or attributes to form connections and associate with each other. This phenomenon is based on the principle that individuals in a network are more likely to connect with others who share similar traits, interests, or demographics (McPherson et al., 2001).
To test whether LLM agents exhibit homophily, we perform the following experiment: We generate nodes with randomly generated attributes regarding a hobby (randomly chosen among three hobbies), a favorite color (randomly chosen among three colors), and a location within the US (randomly chosen among three US locations) and provide the attributes of the other nodes and the node’s own attributes, and each node is tasked to form up to $\delta=5$ links with others. For each node $i$, we provide it with the features $x_{j}$ of all non-neighbors $j$ of $i$. The seed network is taken to be the empty graph.
We run ten simulations for networks with $n=50$ nodes and $\delta=5$. We compare with the Random null model, which chooses $\delta=5$ links at random.
To measure the presence of homophily, we measure the attribute assortativity coefficient of each of the three features (hobby, color, location). For each property $P$ which takes $K$ distinct values $P_{1},\dots,P_{K}$ (indexed by $k$ or $l$), its assortativity coefficient $R$ is defined as
$R=\frac{\sum_{k=1}^{K}M_{kk}-\sum_{k=1}^{K}a_{k}b_{k}}{1-\sum_{k=1}^{K}a_{k}b_{k}}.$
Here $M$ represents the mixing matrix. Its elements $M_{kl}$ reflect the proportion of edges connecting two nodes with values $P_{k}$ and $P_{l}$, respectively. We define $a_{k}=\sum_{l=1}^{K}M_{kl}$ and $b_{k}=\sum_{l=1}^{K}M_{lk}$. Assortativity ranges from $-1$ to $+1$. A positive assortativity indicates nodes preferentially connect to similar ones, forming a homophilous network. Conversely, a negative assortativity suggests connections primarily occur between dissimilar nodes, indicating heterophily.
In Figure 3(a), we observe a consistent pattern of positive assortativity across all temperature ranges, which is higher for lower temperatures. Moreover, the observed assortativity at all temperatures significantly differs from Random, which has assortativity close to 0 ($P<0.001$, t-test).
Moreover, we present the probability of each reason being mentioned in SI Section 3. We find that shared hobbies appear to be much more important than shared location or favorite color. This is consistent with our intuition that, similarly to humans, LLMs care more about hobbies than other superficial attributes when forming social connections. In token, we conclude homophily emerges for the networks generated by LLMs.
Macro-Level Principles
(a) Assortativity Coefficients
(b) Network Instances
(c) Communities Statistics
Figure 3. Results for Principle 3 (homophily). (a) Figure 3(a) shows the assortativity coefficients for each network attribute created based on Principle 3. Each node can form at most $\delta=5$ links with the other nodes based on their attributes. We observe that for all temperatures, we have positive assortativity ($P<0.001$, t-test with 0), which indicates homophily. The assortativities are significantly different for all temperatures from random or heterophilous ($P<0.001$, t-test comparing with Random). (b) Networks created based on Principle 3. Each node can form at most $\delta=5$ links with the other nodes based on their attributes. The edge widths correspond to the number of common attributes shared between the nodes. The node colors correspond to the communities found by running the Louvain algorithm. (c) Louvain modularity, number of communities, and community sizes of networks created based on Principle 3. Lower temperatures yield higher modularity ($P<0.001$, t-test) and fewer and larger communities ($P<0.001$, t-test). For smaller temperatures, the communities are more “well-separated”. The error bars correspond to 95% confidence intervals.
Principle 4: Community Structure
The community structure of networks refers to the organization of nodes or individuals within a network into distinct and densely interconnected groups or clusters (Newman and Girvan, 2004; Newman, 2006; Blondel et al., 2008; Clauset et al., 2004). Identifying community structures is crucial for understanding the overall dynamics of a network, as it reveals patterns of relationships and interactions that might not be apparent at the global level.
Both triadic closure and homophily contribute to the formation of community structures. By examining how these two factors contribute to network formation, we aim to gain insights into the underlying mechanisms driving community dynamics in LLM-generated networks. We employ the simulation results presented in the synthetic networks to determine whether community structure in networks generated by LLMs emerges from triadic closure or homophily.
First, we consider the networks generated in Figure 2. We examine how LLM agents’ choices strengthen the network’s community structure. Specifically, we leverage the fact that the SBM graph has a preexisting community structure and measure how the newly formed links reinforce such a structure. Visual inspection shows that the newly added links, represented by the bold edges, happen mostly within each cluster, reinforcing the community structure. This is further quantitatively verified by the fraction $\hat{p}$ newly created inter-community edges. We find that $\hat{p}$ is significantly higher than 0.5 ($P<0.001$, t-test comparing with 0.5). This indicates that most edges are within the same community, strengthening the community structure.
Next, we investigate the community structure resulting from homophily using modularity maximization (Blondel et al., 2008) (Figure 3). Modularity quantifies the discrepancy between the actual number of edges within communities and the expected number in a random network with identical node count and degree distribution, following the Chung-Lu model (Chung and Lu, 2004). This model presumes that nodes maintain their weighted degree, with edges randomly distributed. The weighted modularity $Q$ (Clauset et al., 2004) for a graph with edge weights $w_{ij}$ and $C$ communities is defined as
$Q=\sum_{c=1}^{C}\left[\frac{L_{c}}{W}-r\left(\frac{k_{c}}{2W}\right)^{2}\right].$
Here $W$ represents the total edge weights, $L_{c}$ the intra-community link weights for community $c$, $k_{c}$ the total weighted degree within community $c$, and $r$ the resolution parameter, set to 1 for our analysis. High modularity values (e.g., greater than 0.5) indicate significant community structuring, diverging from the random model. [^8] Given the NP-Hard nature of maximizing $Q$, we employ the Louvain algorithm (Blondel et al., 2008) to approximate the highest possible modularity.
For the network’s weights, we use the number of common attributes shared between each pair of nodes: $w_{ij}=\left|\left\{k:x_{i}^{(k)}=x_{j}^{(k)}\right\}\right|$ for each link $(i,j)$ in the final network. Here, $x_{i}^{(k)}$ and $x_{j}^{(k)}$ correspond to the $k$-th features of $x_{i}$ and $x_{j}$, respectively.
In Figure 3(b), various colors represent the communities identified by the Louvain algorithm at different temperatures. Notably, communities appear more distinct at lower temperatures, likely due to reduced randomness in decision-making at these temperatures. Additionally, Figure 3(c) presents the distribution of Louvain modularity values across simulations, indicating consistent community structure with positive modularity at all temperatures, confirmed by a t-test against a modularity of $Q=0$ for a random graph ($P<0.001$). Moreover, lower temperatures lead to fewer, larger communities as per the Louvain algorithm’s findings ($P<0.001$, t-test).
Our results demonstrate that community structures manifest in networks generated by LLMs, driven by both triadic closure and homophily.
Principle 5: Small-World
(a) Network Instances for $\beta=0.25$
(b) Network Instances for $\beta=0.5$
(c) Network Instances for $\beta=0.75$
Figure 4. Simulation results for Principle 5 (small world). Network instances for the networks created according to Principle 5 using the altered Watts-Strogatz Model for node count $n=50$, average degree $k=5$, rewriting probability $\beta\in\{0.25,0.5,0.75\}$, together with plots of the average clustering coefficient $C$ and the average shortest path length $L$. The comparison is made with respect to a Watts-Strogatz graph with $n=50,k=5,\beta\in\{0.25,0.5,0.75\}$. The error bars correspond to 95% confidence intervals.
(a) $\beta=0.25$
(b) $\beta=0.5$
(c) $\beta=0.75$
(d) $\beta=0.25$
(e) $\beta=0.5$
(f) $\beta=0.75$
Figure 5. Fitted results for Principle 5 (small world). Top: Regression plots relating average shortest path length ($L$) and average clustering coefficient ($C$) with $n$ for $\beta\in\{0.25,0.5,0.75\}$ and $k=5$. Value $a$ in legends represents the slope of the regression lines. Bottom: Estimated values $\hat{\beta}$ of $\beta\in\{0.25,0.5,0.75\}$ for LLM-generated networks based on matching the average clustering coefficient and difference in the average shortest path between LLM-generated networks and Watts-Strogatz with the estimated rewiring probability $\hat{\beta}$. We report the $P$-values of the t-test comparing the average shortest path length of the LLM-generated networks and the average shortest path length of the Watts-Strogatz graphs with rewiring probability $\hat{\beta}$. *: $P<0.05$; **: $P<0.01$, and ***: $P<0.001$.
The small-world phenomenon is characterized by networks where nodes are interconnected in tight clusters, yet the average distance between any two nodes remains relatively short, typically scaling logarithmically with the network size (Watts and Strogatz, 1998; Kleinberg, 2000). This balance between high clustering and short path lengths characterizes small-world networks.
A small-world network is defined by its average shortest path length $L$, which grows logarithmically with the size of the network $n$, [^9] As per the definition in (Jacobs and Watts, 2021). expressed as
$L\sim\log(n).$
Our analysis utilizes the Watts-Strogatz model (Watts and Strogatz, 1998) as a benchmark to investigate whether LLMs can generate networks exhibiting small-world characteristics. This model has a delicate balance between local clustering and short average path lengths: Nodes tend to form clusters or groups (triadic closure), exhibiting a high level of interconnectedness within these local neighborhoods, whereas at the same time, the existence of a few long-range connections ensures that the entire network is reachable with relatively few steps (Park et al., 2018; Lyu et al., 2022; Jahani et al., 2023).
We employ a modified version of the model, where edge rewiring is informed by LLM queries, based on the current network structure. The generation process is parametrized by the number of nodes ($n$), average degree ($k$), and the rewiring probability ($\beta$). See details in Methods and Materials.
We generated networks of various sizes, ranging from $n=10$ to $n=100$, to explore the relationship between the network size ($n$) and two key metrics: the average shortest path length ($L$) and the average clustering coefficient ($C$). For this analysis, we considered values of $\beta$ set at 0.25, 0.5, and 0.75, with a fixed $k=5$ to serve as a consistent parameter. Visual representations and average clustering coefficients are presented in Figures 4(a), 4(b) and 4(c).
However, when directly compared with the Watts-Strogatz model, the networks generated by the LLMs do not precisely replicate the characteristics of Watts-Strogatz networks for the corresponding the rewiring probabilities ($\beta$). As illustrated in right panels of Figures 4(a), 4(b) and 4(c), we find weak evidence that the LLM-generated networks share the same average shortest path length as the Watts-Strogatz model for the rewiring probabilities ($\beta$) of 0.25, 0.5, and 0.75, with $P<0.1$ from a t-test comparing the average shortest path lengths. Additionally, LLM-generated networks exhibit a larger average clustering coefficient than those of the Watts-Strogatz model for the same rewiring probabilities $\beta$, also with a significance level of $P<0.1$ in the t-test comparisons. This discrepancy suggests the differences in the network structure and connectivity patterns between the LLM-generated networks and the classical Watts-Strogatz model.
We also provide regressions analysis by examining the correlation between the average shortest path length and average clustering coefficient versus $\log(n)$ (refer to Figure 5(a), Figure 5(b), and Figure 5(c)). We found that across all tested temperatures, the relationships were statistically significant, with most regressions yielding $P<0.001$. This indicates that the average shortest path length increases proportionally with $\log(n)$. Similarly, for the average clustering coefficient, we demonstrated that it inversely scales with $1/\log(n)$, with the majority of regression analyses also showing $P<0.001$. These findings align with the small-world properties of organizational networks as documented in the study by (Jacobs and Watts, 2021), suggesting that these characteristics are not only prevalent but also predictable across different network sizes.
To quantify how LLM-generated networks resemble Watts-Strogatz networks, we fit the estimated $\hat{\beta}$ values for each LLM-generated network. [^10] We conducted a binary search to identify the $\hat{\beta}$ values for which the Watts-Strogatz networks’ average clustering coefficients match those of the LLM-generated networks. In Figures 5(d), 5(e) and 5(f), we plot the estimated values for $\hat{\beta}$ for each value of $\beta$ and each temperature. Here $P$-values result from a t-test comparing with the average shortest path length of Watts-Strogatz with rewiring probability $\hat{\beta}$. These results show that while the average shortest path lengths are not identical, they are sufficiently close, with the differences not being statistically significant at the $0.1$ level for most temperature settings.
In conclusion, our analysis demonstrates that LLM-generated networks exhibit key small-world properties, with logarithmic scaling of average shortest path lengths and inverse logarithmic scaling of average clustering coefficients. While these networks do not perfectly align with the Watts-Strogatz model, they exhibit similar structural characteristics.
Decisions on Real-World Networks
To investigate the behavior of LLMs in real-world network formation contexts, we employ the Facebook100 dataset (Traud et al., 2012) and a discrete choice modeling framework (Overgoor et al., 2019; McFadden, 1972). This dataset comprises “friendship” networks from one hundred American colleges and universities, captured at a specific moment in time from Facebook’s online social network. We model the network formation process as a discrete choice process, wherein nodes are sequentially prompted to form connections from a set of available alternatives (see Methods and Materials).
We use datasets that refer to three US colleges – Caltech36, Swarthmore42, and UChicago30. They represent different sizes of networks: $n=769$, $n=1,659$, and $n=6,591$ nodes, respectively. For the Caltech36 ($n=769$) and Swarthmore42 ($n=1,659$) datasets, we consider all of the network nodes to be nodes that have to make a choice. For the UChicago30 dataset, we consider a randomly sampled subset of 2,000 nodes because of the limited context window of GPT models. We set the size of the alternative sets $A_{t}$ to 20, and thus, the accuracy of random guessing is 5%.
Figure 6. Regression results for real-world network experiments. Multinomial logit coefficients for three networks from the Facebook100 dataset. The standard errors of the estimates are shown with error bars. The full regression tables can be found in SI Section 1. All coefficients are statistically significant, where the null hypothesis corresponds to the respective parameter being equal to 0.
We regress network formation decisions on standardized scores reflecting the three micro-level principles. We present the regression results in Figure 6.111111More detailed results can be found in Table S1 in SI Section 1. Our analysis reveals that the degree, the number of common neighbors, and the number of shared interests are all statistically significant factors influencing the choices of LLMs ($P<0.001$) across all datasets. Furthermore, we find that LLMs place the greatest emphasis on the number of common attributes ($\hat{\theta}_{\mathrm{H}}$), followed by the number of common neighbors ($\hat{\theta}_{\mathrm{TC}}$), and then the degree ($\hat{\theta}_{\mathrm{PA}}$).
These results suggest that LLMs exhibit a strong preference for homophily, followed by a tendency towards triadic closure, and finally, preferential attachment. Specifically, the relationship $\hat{\theta}_{\mathrm{H}}>\hat{\theta}_{\mathrm{TC}}>\hat{\theta}_{\mathrm{PA}}$ holds across all examined datasets and temperature settings.
Further insights are provided in SI Section 4, where we provide the reasoning analysis. The analysis points to a marked emphasis on homophily, indicating that nodes with similar characteristics—such as student/faculty status, high school, and graduation year—are more likely to connect. In contrast, triadic closure and preferential attachment have a relatively lesser impact.
Our study also includes a comparison of LLMs’ predictive capabilities against random guessing in a link prediction framework. We assess the percentage increase in accuracy, defining accuracy as the probability of selecting an existing neighbor from a set of alternatives, as opposed to a random selection from 20 possible options. As detailed in SI Section 1, LLMs significantly outperform random guessing in accuracy.
Furthermore, we investigate whether LLMs reinforce community structure post-network formation. We run the Louvain algorithm 10 times for both the initial graph $G_{1}$ and the final graph $G_{T}$ and perform a t-test between the resulting modularity scores (see SI Section 1). We observe that the modularity increases in all cases and is statistically significant. This suggests that LLMs make choices that naturally yield a reinforcement of the community structure.
Additionally, to explore the “small-world” phenomenon, we measure the percent change in average shortest path length between the pre- and post-decision networks (detailed in SI Section 1). Our findings reveal a minor decrease in average path length in most cases, suggesting a slight strengthening of the small-world phenomenon.
In summary, our analysis demonstrates that while homophily, triadic closure, and preferential attachment are integral to the network formation behaviors of LLMs, homophily is the dominant factor. Furthermore, both community structure and the small-world phenomenon are reinforced after network formation decisions of LLMs.
Discussion
Our study conducted a comprehensive evaluation of LLMs’ network formation preferences, encompassing both micro-level network principles such as preferential attachment, triadic closure, and homophily and macro-level network principles like community structure and the small-world phenomenon. Our findings indicate that networks generated by multiple LLMs do exhibit these properties, particularly when primed with network statistics such as the number of mutual friends or the degree of alternatives. Furthermore, through discrete choice modeling, we explored the emergence of these properties in real-world networks. Our real-world network results reveal that the agents’ selections are predominantly driven by homophily, followed by triadic closure, and then preferential attachment. Additionally, we observed a strengthening of community structure in real-world network experiments.
Our study enhances our understanding of the behavior of multi-LLMs. For example, our findings reveal different strengths of network formation properties, suggesting that when LLMs are employed to coordinate social networks in social or work environments, they may exhibit behaviors similar to humans. This insight contributes to developing socially aware LLMs that align with human preferences, with potential applications ranging from chatbots and virtual assistants to collaborative systems and social media platforms.
Furthermore, our study may pave the way for innovative applications of LLMs in network science and computational social science research. Given that LLMs adhere to fundamental network principles and exhibit behaviors akin to human network formation, they hold significant potential for use in simulations of social systems and agent-based modeling. As LLM agents gain increasing ability to exhibit more complex and nuanced behaviors, such as negotiation, collaboration, and competition, this presents opportunities to replicate networked environments like marketplaces or organizational structures. In these settings, LLMs could effectively model decision-making processes and interaction dynamics.
Looking ahead, we propose three promising directions for future research. First, we propose investigating LLM behavior in more complex interactions, such as simulated dialogs or specific environments like companies, schools, or parties. By examining LLMs in these contexts, we can better understand their network formation preferences and how they adapt to different social dynamics. This research could shed light on the nuances of LLM behavior in realistic settings and inform their integration into diverse social environments. Second, we suggest exploring the application of our findings in specific settings, particularly in organizational contexts. For example, considering the potential use of LLMs to assist HR professionals, our study’s insights could guide the alignment of LLMs with organizational goals. By leveraging network information, LLMs could help HR departments identify their firms’ best fit or talent, optimizing recruitment and talent management processes. Finally, we propose using our methods to create realistic synthetic networks. These synthetic datasets can serve as benchmarks for evaluating graph learning methods, thus addressing the scarcity of existing graph benchmarks. By adjusting parameters like temperature, we can generate diverse networks to test graph neural network performance under various conditions. In addition, our approach can generate artificial data resembling real-world data while adhering to privacy regulations. This has the potential to expedite research in graph learning and trustworthy machine learning by providing realistic network datasets generated by LLMs.
1. Methods
Methods and Materials
# Task Your task is to select a set of people to be friends with. # Profile Your profile is given below after chevrons: <PROFILE>$F(\{i_{t}\})$</PROFILE> # Candidate Profiles The candidate profiles to be friends with are given below after chevrons: <PROFILES>$F(A_{t})$</PROFILES> # Output The output should be given a list of JSON objects with the following structure [ {{ "name" : name of the person you selected, "reason" : reason for selecting the person }}, ... ] # Notes - The output must be a list of JSON objects ranked in the order of preference. - You can make at most $\delta$ selections.
ALGORITHM 1 Prompt used to implement $\mathcal{Q}(A_{t},i_{t},\delta)$.
Experimental Procedure
In our study, we performed experiments to assess whether key network principles at both the micro-level (such as preferential attachment, triadic closure, and homophily) and the macro-level (including community structure and weak ties) align with classical network models. Subsequently, we utilized real-world networks to determine the factors that are most heavily weighted by LLMs.
Our experiments span a time series of $T$ steps, with a sequence of network structures denoted as $G_{1},G_{2},\dots,G_{T}$ with vertex sets $V_{1},\dots,V_{T}$. The initial network, $G_{1}$, is referred to as the seed network. At each step $t$, we select a query node $i_{t}$ (which may either be a new arrival or an existing node in the graph) and assign it the task of forming new links. This is accomplished by selecting nodes from a set of alternatives $A_{t}$ (meaning potential candidates for link formation) and initiating a query call $\mathcal{Q}(A_{t},i_{t},\delta)$ to the LLM (as outlined in Algorithm 1) to create up to $\delta$ new links. The edge set selection process involves presenting the LLMs with personal or network features of the alternatives, denoted as $F(A_{t})=\left\{f_{a}:a\in A_{t}\right\}$, which may include information such as the neighbors of the nodes, node degrees, common connections with $i_{t}$, and community memberships, formatted in JSON. We adopt a zero-shot learning approach, avoiding the provision of examples to the model to prevent bias, in line with relevant studies such as (Brookins and DeBacker, 2023). This approach allows for the exploration of the innate preferences of LLMs.
We employ multiple temperatures to account for the variability in response generation by LLM systems, which is also observed in classical statistical models of network formation (Jackson, 2008). Our study conducts experiments using three temperatures: 0.5, 1.0, and 1.5.
Moreover, the model is tasked with outputting a JSON object indicating the node chosen for link formation and the rationale behind the choice. This approach is adopted because LLMs have demonstrated proficiency in processing code-like structures, such as HTML and JSON.
Summarizing the Reasoning of the LLM Agents
In addition to analyzing the innate preferences of LLM agents, we employ an interview-like approach to probe the rationale behind their decisions. Following the LLMs’ decisions on network formation, we inquire about their reasoning for making those choices. We then aggregate all the provided rationales using GPT models (GPT-3.5 for the synthetic networks and GPT-4 for the real-world networks) by asking it to identify the common themes, resulting in the creation of multiple categories. To determine the category to which each LLM’s reasoning belongs, we subsequently ask the GPT models to classify each rationale into the appropriate category.
Details for Small-World Experiments
The algorithm of generating LLM networks is described as follows:
  1. (1)
    Similarly to Watts-Strogatz, we first create a ring network with $n$ nodes. After that, for each node $[n]$, we create $k$ edges where $k/2$ edges connect to its rightmost neighbors and $k/2$ edges connect to its leftmost neighbors.
  2. (2)
    To create $G_{t}$, for each node $[n]$, we take its $k/2$ rightmost neighbors and rewire them with probability $\beta$. For each of the $k/2$ rightmost neighbors that are to be rewired, we make one query to the LLM, which indicates how the edge will be rewired. The choice is made by providing the LLM with all the network nodes and each node’s neighbors (i.e., the network structure).
The model closely resembles the Watts-Strogatz model, with the primary distinction being the method of edge rewiring. Instead of randomly selecting edges for rewiring, as in the Watts-Strogatz model, we determine the rewiring of an edge by inquiring about the LLM and providing it with the current network structure.
The Discrete Choice Model in Real-World Network Experiments
For each node $i_{t}$ that we consider at time $t$, we randomly remove one of its current friends from the real-world network. After we remove a neighbor for each of $i_{1},\dots,i_{T}$, we end up with the network $G_{1}$, which we use as a seed network for the LLM agents.
Subsequently, during the link formation process, we present each node $i_{t}$ with a set of candidate nodes (denoted by $A_{t}$), comprising one of the previously removed friends and other nodes that are not their friends. We then instruct the LLM to form a link with one of the candidates, providing the attributes of the candidates and the social network structure to aid its decision-making. These choices are made sequentially.
We use the utility of the model for each node for each sequential decision of network formation:
$U_{ij,t}=\theta_{\mathrm{PA}}\log d_{j,t}+\theta_{\mathrm{H}}\log w_{ij}+\theta_{\mathrm{TC}}\log c_{ij,t}+\epsilon_{ij,t}.$
In this equation, $\theta_{\mathrm{PA}}$ measures the strength of preferential attachment based on the degree $d_{j,t}$ of $j$ at step $t$, $\theta_{\mathrm{H}}$ measures the strength of homophily based on the similarity $w_{ij}$ (i.e. number of common attributes) between $i$ and $j$, and $\theta_{\mathrm{TC}}$ measures the strength of triadic closure, based on the number of common neighbors $c_{ij,t}$ between $i$ and $j$ at step $t$. The error term $\epsilon_{ij,t}$ is distributed as i.i.d. standard Gumbel. [^12] The standard Gumbel distribution has CDF $e^{e^{-x}}$. All variables are first normalized based on their range, and then the log transformation is taken.
The multinomial logit model (MNL) indicates that the probability that $i$ links to $j$ at step $t$ is given by
$p_{ij,t}=\Pr\left[\mathrm{argmax}_{r\in A_{t}}U_{ir,t}=j\right]=\frac{d_{j,t}^{\theta_{\mathrm{PA}}}w_{ij}^{\theta_{\mathrm{H}}}c_{ij,t}^{\theta_{\mathrm{TC}}}}{\sum_{r\in A_{t}}d_{r,t}^{\theta_{\mathrm{PA}}}w_{ir}^{\theta_{\mathrm{H}}}c_{ir,t}^{\theta_{\mathrm{TC}}}}.$
Given a sequence of nodes $i_{1},\dots,i_{T}\in V$ and choices (denoted by subscripted $j$) $j_{1}\in A_{1},\dots,j_{T}\in A_{T}$, the parameters can be found by maximizing the log-likelihood function. To get the standard errors of the coefficients and the corresponding $P$-values, we follow the process outlined in (Overgoor et al., 2019).
Estimating the Parameters of the Discrete Choice Model
To estimate the parameters of the discrete choice model, we optimize the following log-likelihood function
$(\hat{\theta}_{\mathrm{PA}},\hat{\theta}_{\mathrm{TC}},\hat{\theta}_{\mathrm{H}})=\operatorname*{arg\,max}_{(\theta_{\mathrm{PA}},\theta_{\mathrm{TC}},\theta_{\mathrm{H}})\in\mathbf{R}^{3}}\sum_{t=1}^{T}\bigg{(}\theta_{\mathrm{PA}}\log d_{j_{t},t}+\theta_{\mathrm{H}}\log w_{i_{t}j_{t}}+\theta_{\mathrm{TC}}\log c_{i_{t}j_{t},t}-\log\big{(}\sum_{r\in A_{t}}d_{r,t}^{\theta_{\mathrm{PA}}}w_{i_{t}r}^{\theta_{\mathrm{H}}}c_{i_{t}r,t}^{\theta_{\mathrm{TC}}}\big{)}\bigg{)},$
where $i_{1},\dots,i_{T}$ are the chooser nodes (i.e., the LLM agents who want to form a link), and $j_{1},\dots,j_{T}$ are the nodes which are chosen from the alternative sets $A_{1},\dots,A_{T}$. The likelihood function is convex, and we optimize it with the L-BFGS-B method (Liu and Nocedal, 1989). The standard errors of the coefficients are approximated as $\sqrt{-H^{-1}}$ where $H$ is the Hessian matrix of the log-likelihood at $(\hat{\theta}_{\mathrm{PA}},\hat{\theta}_{\mathrm{TC}},\hat{\theta}_{\mathrm{H}})$ (cf. (Overgoor et al., 2019; Train, 2009)).
Data and Code Availability
Data and code are available at GitHub: https://github.com/papachristoumarios/llm-network-formation.
Acknowledgements.
The authors would like to thank Jon Kleinberg, Yanbang Wang, Yuanqi Du, the attendees of the Learning on Graphs NYC meetup, and the participants of Cornell AI, Policy, and Practice working group for their valuable feedback on the current version of the paper. M.P. is supported by a scholarship from the Onassis Foundation (Scholarship ID: F ZT 056-1/2023-2024).
Appendix A Regression Table
In Table S.1, we report the regression coefficient for the regression in the real-world network data. The first column corresponds to the temperature, the next three columns correspond to the fitted coefficients from the regression model of Section 1.C (also shown in Figure 5) accompanied by the standard errors (in parentheses) and the $P$-values indicated by stars (the null hypothesis corresponds to the parameters being set to 0). Next, LL corresponds to the log-likelihood of the fitted model, and AIC corresponds to the Akaike Information Criterion. Finally, we report the percent change in the accuracy compared to random guessing, the percent change in the average path length (as a measure of the small-world phenomenon), and the clustering coefficient (as a measure of the small-world phenomenon and the triadic closure), as well as the $t$-statistic for the change in modularity ($Q$) between the ground truth network dataset (before the edge deletions) and the network after the network formation process.
We observe that $\hat{\theta}_{\mathrm{H}}>\hat{\theta}_{\mathrm{TC}}>\hat{\theta}_{\mathrm{PA}}>0$ accross all settings. LLM agents do better than random guessing, reinforce the small-world phenomenon, and weaken the triadic closure. Finally, the community structure is strengthened after new links are formed.
Temp. $\hat{\theta}_{\mathrm{PA}}$ $\hat{\theta}_{\mathrm{H}}$ $\hat{\theta}_{\mathrm{TC}}$ LL AIC % Change % Change % Change $\Delta Q$ (t-stat)
Acc. $L$ $C$
Caltech36 (769 nodes, 33,312 edges)
0.5 0.41*** (0.31) 1.95*** (0.64) 0.59*** (0.37) -1,377.47 2,762.94 171.8 -0.008 -9.94 3.45**
1.0 0.36*** (0.13) 1.85*** (0.52) 0.58*** (0.33) -1,435.07 2,878.13 179.6 -0.18 -11.08 3.49**
1.5 0.36*** (0.16) 1.72*** (0.31) 0.55*** (0.20) -1,522.47 3,052.94 127.6 -0.06 -11.46 3.37**
Swarthmore42 (1,659 nodes, 12,2100 edges)
0.5 0.18*** (0.13) 1.62*** (0.23) 0.65*** (0.08) -2,838.33 5,684.66 124.2 0.01 -11.46 7.42***
1.0 0.26*** (0.07) 1.70*** (0.35) 0.58*** (0.12) -2,927.99 5,863.97 91.6 -0.10 -4.25 1.96*
1.5 0.19*** (0.14) 1.50*** (0.33) 0.59*** (0.10) -3,139.42 6,286.83 87.39 -0.20 -4.52 4.03***
UChicago30 (6,591 nodes, 416,206 edges)
0.5 0.23*** (0.06) 2.00*** (0.24) 0.41*** (0.09) -3,444.33 6,896.67 217.2 -0.24 -2.52 7.46*** [0.34]
1.0 0.23*** (0.10) 1.98*** (0.17) 0.38*** (0.06) -3,578.18 7,164.36 219.2 -0.12 -2.66 9.56*** [1.05]
1.5 0.22*** (0.12) 1.78*** (0.27) 0.41*** (0.07) -2,033.49 4,074.98 222.4 -0.17 -2.42 10.19*** [0.24]
Notes $\hat{\theta}_{\mathrm{PA}}$ = Coefficient of log degree, $\hat{\theta}_{\mathrm{H}}$ = Coefficient of log # of common attributes, $\hat{\theta}_{\mathrm{TC}}$ = Coefficient of log # common neighbors
LL = Log-likelihood, AIC = Akaike Information Criterion
Acc. = Accuracy, $L$ = Average Path Length, $C$ = Average Clustering Coefficient, $\Delta Q$ (t-stat) = Modularity change t-statistic
${}^{*}:P<0.05,\;^{**}:P<0.01,\;^{***}:P<0.001$
Table S.1. Multinomial logit coefficients for three networks from the Facebook100 dataset. The standard errors of the estimates are shown in parentheses. The null hypothesis corresponds to the respective parameter being equal to 0. We report the percent change in accuracy, average path length, and average clustering coefficient compared to the initial network (before the deletion of edges). For the change in modularity, we run the Louvain algorithm ten times and perform a t-test with the resulting modularities. For the UChicago30 dataset, we report the t-statistic value in the subgraph induced by the 2,000 sampled nodes, since the newly added edges would have a very small effect on the change in the community structure if we were to measure it in the whole network. We also report the modularity change (t-statistic) of the whole graph inside brackets.
Appendix B Analytical Regression Tables
We present the regression tables for all the combinations of coefficients for each of the three real-world network datasets. For all datasets, we observe that $\hat{\theta}_{\mathrm{PA}}$ is smaller than both $\hat{\theta}_{\mathrm{TC}}$ and $\hat{\theta}_{\mathrm{H}}$ in all models where any pair is included. Similarly, $\hat{\theta}_{\mathrm{TC}}$ is always smaller than $\hat{\theta}_{\mathrm{H}}$ in all models that are both included. Finally, note that whenever only $\hat{\theta}_{\mathrm{PA}}$ and $\hat{\theta}_{\mathrm{TC}}$ are considered, then $\hat{\theta}_{\mathrm{PA}}<0$ and the result is not statistically significant ($P>0.05$).
Temp. $\hat{\theta}_{\mathrm{PA}}$ $\hat{\theta}_{\mathrm{H}}$ $\hat{\theta}_{\mathrm{TC}}$ Log Likelihood AIC
0.5 0.50*** (0.04) -2,236.36 4,476.71
0.5 2.78*** (0.07) -1,511.42 3,026.85
0.5 1.53*** (0.04) -1,506.01 3,016.02
0.5 0.64*** (0.06) 2.99*** (0.19) -1,414.71 2,835.41
0.5 -0.02 (0.10) 1.53*** (0.08) -1,505.95 3,017.90
0.5 1.43*** (0.11) 0.82*** (0.06) -1,406.39 2,818.78
0.5 0.41*** (0.31) 1.95*** (0.64) 0.59*** (0.37) -1,377.47 2,762.94
1.0 0.48*** (0.04) -2,242.85 4,489.70
1.0 2.69*** (0.07) -1,558.67 3,121.34
1.0 1.47*** (0.04) -1,556.68 3,117.37
1.0 0.58*** (0.05) 2.86*** (0.08) -1,473.13 2,952.26
1.0 -0.04 (0.06) 1.47*** (0.06) -1,556.24 3,118.48
1.0 1.40*** (0.28) 0.79*** (0.06) -1,457.76 2,921.52
1.0 0.36*** (0.13) 1.85*** (0.52) 0.58*** (0.33) -1,435.07 2,878.13
1.5 0.50*** (0.04) -2,233.25 4,470.50
1.5 2.51*** (0.07) -1,646.83 3,297.65
1.5 1.36*** (0.04) -1,636.20 3,276.40
1.5 0.57*** (0.07) 2.65*** (0.13) -1,559.57 3,125.15
1.5 -0.01 (0.05) 1.37*** (0.04) -1,636.19 3,278.37
1.5 1.29*** (0.11) 0.75*** (0.06) -1,546.78 3,099.55
1.5 0.36*** (0.16) 1.72*** (0.31) 0.55*** (0.20) -1,522.47 3,052.94
Note ${}^{*}:P<0.05,\;^{**}:P<0.01,\;^{***}:P<0.001$
Table S.2. Multinomial logit coefficients for Caltech36. The standard errors of the estimates are shown in parentheses.
Temp. $\hat{\theta}_{\mathrm{PA}}$ $\hat{\theta}_{\mathrm{H}}$ $\hat{\theta}_{\mathrm{TC}}$ Log Likelihood AIC
0.5 0.33*** (0.03) -4,978.28 9,960.56
0.5 2.91*** (0.05) -3,027.03 6,058.06
0.5 1.37*** (0.02) -3,014.28 6,032.57
0.5 0.44*** (0.15) 2.99*** (0.10) -2,948.91 5,903.82
0.5 -0.18*** (0.09) 1.36*** (0.07) -3,002.17 6,010.35
0.5 1.42*** (0.11) 0.72*** (0.07) -2,847.16 5,700.32
0.5 0.18*** (0.13) 1.62*** (0.23) 0.65*** (0.08) -2,838.33 5,684.66
1.0 0.38*** (0.03) -4,959.02 9,922.04
1.0 2.83*** (0.05) -3,119.06 6,242.11
1.0 1.32*** (0.02) -3,118.13 6,240.26
1.0 0.50*** (0.16) 2.92*** (0.09) -3,018.85 6,043.71
1.0 -0.11** (0.09) 1.32*** (0.07) -3,113.21 6,232.43
1.0 1.41*** (0.15) 0.69*** (0.11) -2,947.89 5,901.78
1.0 0.26*** (0.07) 1.70*** (0.35) 0.58*** (0.12) -2,927.99 5,863.97
1.5 0.36*** (0.03) -4,952.51 9,909.02
1.5 2.64*** (0.05) -3,324.71 6,653.43
1.5 1.24*** (0.02) -3,306.06 6,616.11
1.5 0.44*** (0.11) 2.71*** (0.05) -3,241.67 6,489.35
1.5 -0.14*** (0.10) 1.24*** (0.05) -3,298.07 6,602.13
1.5 1.30*** (0.14) 0.67*** (0.09) -3,150.29 6,306.57
1.5 0.19*** (0.14) 1.50*** (0.33) 0.59*** (0.10) -3,139.42 6,286.83
Note ${}^{*}:P<0.05,\;^{**}:P<0.01,\;^{***}:P<0.001$
Table S.3. Multinomial logit coefficients for Swarthmore42. The standard errors of the estimates are shown in parentheses.
Temp. $\hat{\theta}_{\mathrm{PA}}$ $\hat{\theta}_{\mathrm{H}}$ $\hat{\theta}_{\mathrm{TC}}$ Log Likelihood AIC
0.5 0.28*** (0.02) -5,983.56 11,971.13
0.5 2.93*** (0.05) -3,637.18 7,278.37
0.5 1.11*** (0.02) -3,740.78 7,485.55
0.5 0.38*** (0.04) 3.06*** (0.21) -3,523.06 7,052.12
0.5 -0.04 (0.08) 1.11*** (0.04) -3,739.31 7,484.61
0.5 1.68*** (0.13) 0.51*** (0.07) -3,477.37 6,960.75
0.5 0.23*** (0.06) 2.00*** (0.24) 0.41*** (0.09) -3,444.33 6,896.67
1.0 0.28*** (0.02) -5,982.66 11,969.32
1.0 2.85*** (0.05) -3,759.78 7,523.56
1.0 1.07*** (0.02) -3,879.04 7,762.08
1.0 0.36*** (0.07) 2.96*** (0.31) -3,649.84 7,305.68
1.0 -0.04 (0.06) 1.07*** (0.03) -3,877.83 7,761.66
1.0 1.67*** (0.11) 0.49*** (0.08) -3,611.48 7,228.95
1.0 0.23*** (0.10) 1.98*** (0.17) 0.38*** (0.06) -3,578.18 7,164.36
1.5 0.30*** (0.03) -3,241.67 6,487.34
1.5 2.71*** (0.06) -2,145.02 4,294.03
1.5 1.03*** (0.02) -2,175.32 4,354.64
1.5 0.37*** (0.08) 2.81*** (0.10) -2,080.67 4,167.34
1.5 -0.01 (0.04) 1.03*** (0.04) -2,175.25 4,356.49
1.5 1.50*** (0.10) 0.51*** (0.04) -2,051.61 4,109.23
1.5 0.22*** (0.12) 1.78*** (0.27) 0.41*** (0.07) -2,033.49 4,074.98
Note ${}^{*}:P<0.05,\;^{**}:P<0.01,\;^{***}:P<0.001$
Table S.4. Multinomial logit coefficients for UChicago30. The standard errors of the estimates are shown in parentheses.
Appendix C LLM Reasoning Summarization
In addition to studying LLM agents’ innate preferences, we use an interview-style approach to understand why they make certain decisions. After LLMs form networks, we ask them about their reasoning. We then use GPT models (GPT-3.5 for Section 1.A. and GPT-4 for Section 1.C) to group their explanations into common themes, creating various categories. To classify each LLM’s reasoning, we have the GPT models assign them to the relevant category.
C.1. Micro-level Principles
In Figure S.7, we show the reasoning of the LLM agents from the networks of Section 1.A.
We observe that for Figure 7(a), the well-connectedness reason becomes less important as the temperature increases while having more mutual friends becomes more important as the temperature increases. This means that the agents emphasize triadic closure in higher temperatures and decrease emphasis on preferential attachment.
For Figure 7(b), we observe an opposite pattern in the LLM reasoning.
A possible explanation for this asymmetry is the fact that different types of features are provided, i.e., in the case of Figure 7(a) the neighbors of each node are provided, whereas in the case of Figure 7(b) the common neighbors with the query node $i_{t}$ are provided.
Finally, Figure 7(c) contains the reasons for Figure 3 based on having a common hobby, a common color, or a common location. We observe that the LLM agents make choices primarily based on sharing the same hobby and then based on a common location or a common color (on an almost equal basis).
(a) Reasons for Figure 1(a)
(b) Reasons for Figure 2(a)
(c) Reasons for Figure 3(b)
Figure S.7. Distribution of reasoning across categories for the networks of Figures 1(a), 2(a), and 3. The classification uses GPT-3.5 on 20 randomly sampled reasons repeated over five times (micro-averaged frequencies). The error bars correspond to 95% confidence intervals.
C.2. Real-World Data
We also present the LLM reasoning for real-world networks. The reasons have been set so that there is a one-to-one correspondence with the regression coefficients. Specifically, the Mutual Friends reason corresponds to $\hat{\theta}_{\mathrm{TC}}$, the Number of friends reason to $\hat{\theta}_{\mathrm{PA}}$, and the Similar attributes reason corresponds to $\hat{\theta}_{\mathrm{H}}$ respectively. Consistently with the results reported in Table S.1, the agents’ reasoning is more geared towards homophily (same attributes), then triadic closure (mutual friends), and finally preferential attachment (number of friends). These results are consistent with the regression.
Figure S.8. Distribution of reasoning across categories for the real-world networks. The classification is done using GPT-4 on a set of 20 randomly sampled reasons repeated over 5 times. In agreement with Table S.1, the multi-LLMs put more weight on homophily, then on triadic closure, and then on preferential attachment. The error bars correspond to 95% confidence intervals.
Appendix D Network Evolution
Here we depict the evolution of the networks generated by the LLM agents.
D.1. Principle 1: Preferential Attachment
We plot the evolution of the LLM-based preferential attachment networks at three timesteps, together with the degree distribution alongside the degree distribution of a BA graph with the same number of nodes. We observe that for the temperature being 0.5 we have a core-periphery like formation which diverges from the BA model, whereas for the temperature being 1.5 the network has the same degree distribution as the BA model.
Figure S.9. Dynamic evolution of networks created based on Principle 1.
D.2. Principle 2: Triadic Closure
We plot the evolution of the LLM-generated networks based on the triadic closure principle, together with the transitivity measure and the algebraic connectivity (which corresponds to the second-smallest eigenvalue of the graph Laplacian). We observe that the algebraic connectivity gradually increases as new edges between the clusters are created. Specifically, the algebraic connectivity reaches a higher value for higher temperatures, indicating the more frequent creation of new intra-cluster edges. Moreover, we observe that the transitivity initially increases and then decreases until it reaches its final value.
Figure S.10. Dynamic evolution of networks created based on Principle 2.
Appendix E Feature Representations for Prompts
Below, we give examples of the features used in the prompt presented in Algorithm 1. The features are formatted as a list of JSON objects which are provided to the prompt.
E.1. Principle 1: Preferential Attachment
For the structure-based information (Figure 1(a)) we have the following features:
"name" : 0,
"neighbors" : [5, 7, 1, 6]
And for the degree-based information (Figure 1(b)) we have the following features:
"name" : 0,
"degree" : 4
E.2. Principle 2: Triadic Closure
For the structure-based information (Figure 2(a)) we have the following features:
"name" : 0,
"common_neighbors" : [5, 7, 1, 6]
And for the degree-based information (Figure 2(b)) we have the following features:
"name" : 0,
"num_common_neighbors" : 4
E.3. Principle 3: Homophily
We have the following features:
"name" : 0,
"favorite_color" : "red",
"hobby" : "hiking",
"location" : "Boston"
E.4. Principle 5: Small-World
We have the following features:
"name" : 0,
"neighbors" : [5, 7, 1, 6]
E.5. Real-World Data
We have the following features:
"name" : 0,
"status" : "student",
"major" : 10,
"second major" : 93,
"accommodation" : "house",
"high_school" : 5,
"graduation_year" : 2008
We note that the initial Facebook100 dataset included gender information as a feature. We chose not to include the gender as one of the features as it has been shown that language models exhibit gender bias (Vig et al., 2020; Kotek et al., 2023; Bordia and Bowman, 2019).